# Persistent Homology-Based Indicator of Orientational Ordering in Two-Dimensional Quasi-Particle Systems applied to Skyrmion Lattices

### v2

@@ Abstract
Two-dimensional (2D) quasi-particle systems, such as magnetic skyrmions, exhibit topological phase transitions between unique 2D phases. However, to identify the phases, an easy and computationally low-cost methodology for capturing the configurational properties of lattice ordering and constructing such an appropriate descriptor that can be easily calculated is, though, not obvious. Here, we propose an indicator for topological phase transitions using persistent homology (PH). PH offers novel insights beyond conventional indicators by capturing topological features derived from the configurational properties of the lattice. The proposed persistent-homology-based topological indicator, which selectively counts stable features in a persistence diagram, effectively traces the lattice’s topological changes, as confirmed by comparisons with the conventionally used measure of the ordering (the magnitude of the orientational order parameter $\langle|\Psi_{6}|\rangle$), typically used to identify lattice phases. We demonstrate the applicability of our indicator to experimental data, showing that it yields results consistent with those of simulations. This experimental validation highlights the robustness of the proposed method for real physical systems beyond idealized simulated systems. While our method is demonstrated in the context of skyrmion lattice systems, the approach is general and can be extended to other two-dimensional systems composed of interacting quasi-particles. As a key advantage, our indicator offers lower computational complexity than the conventionally used measures.
1  Introduction
@@
From a fundamental physics perspective, ensembles of skyrmions, known as skyrmion lattices, can exhibit complex phase behavior, including two-dimensional melting transitions [10]. These transitions are theoretically captured by the Kosterlitz–Thouless–Halperin–Nelson–Young
@@
(KTHNY) framework, which describes the emergence of intermediate hexatic phases and the role of topological defects such as dislocations and disclinations [10]. Although this behavior has been extensively studied in colloidal systems, skyrmion lattices offer distinct advantages due to their tunability and controllable dynamics that can be fully captured in real time and real space, making them an ideal platform for investigating two-dimensional phase transitions [11].
@@
Despite intensive research, a computationally easy methodology for describing the configurational properties of lattices remains elusive. Conventional methods (e.g., using the local orientational order parameter $\psi_{6}(r)$ [5, 11, 12, 13]) are limited by their reliance on ensemble-averaged quantities and are computationally expensive. That is because they normally average the quantities encoded to each particle over all the particles to identify the state in the systems.
@@
To address this, we introduce a framework based on Topological Data Analysis (TDA), a concept from algebraic topology used to analyze the geometric structure of objects, specifically applying Persistent Homology (PH) to extract configurational properties of lattice configurations [14, 15, 16, 17, 18, 19]. The PH provides a multi-scale view of topological features by analyzing how connected components and loops emerge and disappear across a filtration. This approach has seen successful applications in biological and materials systems [20, 21, 22, 23, 24]. PH can capture and describe the microscopic processes involved in the phase transitions [16]. In this work, we adopt a particle-based approach, modeling skyrmions as interacting quasi-particles and focusing on their positional configuration that does not require knowledge of the full spin texture. This abstraction enables an efficient computational treatment and highlights the essential geometric features responsible for phase behavior [3, 4, 25]. To extract physically meaningful insights from the persistence diagrams, we propose a topological indicator [26, 27], the Persistent Generator Count with Relative Stability (PGCRS), which selectively counts only the robust topological features in each homology dimension. PGCRS reliably detect phase transitions in the lattice, correlates strongly with conventional measures, and offers a significantly reduced computational complexity. A distinctive aspect of our approach is inverse analysis [28], to trace persistent generators back to specific real-space configurations, enabling a direct physical interpretation of the microscopic structures responsible for topological phase behavior. This conceptual shift from averaging local order to counting persistent topological features offers a new perspective for understanding 2D phase transitions. In addition to numerical simulations, we apply our framework to experimentally acquired skyrmion configurations, demonstrating the practical applicability of our method for real physical systems. While the focus of this work is on skyrmion systems, the underlying method can be extended to other two-dimensional quasi-particle systems.
This paper is organized as follows: we first describe the simulation setup, experimental procedure and analysis framework. We then present detailed results, including the persistence diagrams, inverse structural analysis, and PGCRS, and conclude with a discussion of implications and potential extensions.
2  Methods
@@
In this work, we treat skyrmions as quasi-particles where the effective inter-particle interactions reflect those of real skyrmion systems, without considering their detailed spin textures [3, 25, 29]. This abstraction allows us to focus on the configurational and topological properties relevant to phase transitions, while significantly simplifying the computational analysis.
The skyrmion lattice states used for persistent homology analysis are generated via Brownian dynamics simulation, specifically a Thiele model-based approach, using the HOOMD-blue package [30]. In this work, the Magnus term is assumed negligible under the high-density skyrmion condition, where the resulting short mean free path limits its influence, and is therefore omitted [31]. Consequently, the Thiele model simulation effectively reduces to a standard Brownian dynamics simulation in the overdamped limit [31]. This simulation models a two-dimensional system of quasi-particles interacting under conditions that reproduce a range of ordering according to different phases, including liquid, hexatic, and solid states [10, 11].
In addition to the simulation data, we apply our method to experimental skyrmion configurations, allowing for a direct comparison with the numerically simulated configurations in the persistent homology analysis.
In the subsequent sections, we first describe the simulation setup, including the inter-particle potential and system initialization, as well as the experimental procedure. We then introduce the local orientational order parameter, its correlation function, and the averaged absolute value of the local orientational order parameter. Finally, we describe the persistent homology analysis and present the construction of the Persistent Generator Count with Relative Stability, our proposed topological indicator.
2.1  Brownian Dynamics Simulation
The Brownian dynamics simulation is performed by using the HOOMD-blue package [30]. The modelled system is a 2D system of $N=65000$ particles with hexagonal packing under periodic boundary conditions. The governing model is classical Brownian dynamics, in which $\beta\varepsilon=1$ is set, where $\beta=1/k_{B}T=1$ denotes the inverse thermal energy and $\varepsilon$ represents the interaction strength [29]. The interaction potential in this system is described by
$U(r)=\varepsilon({\sigma}/{r})^{n},$ (1)
where $\sigma$ represents the characteristic interaction range, $n$ corresponds to the steepness of the potential, and $r$ is the distance from the referenced particle [3, 5, 29]. In this work, $\sigma=1.0$, and $n=8.0$ are chosen. It is worth noting that two commonly used effective skyrmion-skyrmion interaction models are the power-law decaying function, used in our simulations, and the exponentially decaying function [3, 5, 11]. The simulation time step is chosen as $dt=10^{-4}$. The well-equilibrated initial state is prepared under an equilibration process for a duration of $10^{7}$ time steps to ensure relaxation to a steady state. After equilibration, snapshots are recorded every $10^{5}$ time steps for additional $10^{7}$ time steps, ensuring no correlations in the state between snapshots. To generate the state spanning a wide range of topological phases, 22 density values are scanned from $\phi$ = 1.1000 to 1.7000. The system is initialized with the 65000 quasi-particles under constant thermal activation over different systems in which the topological phases change depending on the value of density, which is defined as the number of particles relative to the size of the simulation box (specifically, the density value is tuned by the size of the simulation box).
2.2  Experimental Data
@@
The experimental skyrmion configuration data used in this work are obtained from magnetic multilayer stacks: Ta(5)/Co20Fe60B20(0.9)/Ta(0.07)/MgO(2)/Ta(5) [3, 11]. Magnetic fields are applied both in-plane (IP) and out-of-plane (OOP) using an Evico Magnetics GmbH Kerr microscope, with the sample maintained at a constant temperature of 333.5 K. Skyrmions are nucleated by applying an IP field pulse under a constant OOP field. The lattice is subsequently equilibrated using an oscillating OOP field at 100 Hz (with amplitudes up to $60\,\upmu\mathrm{T}$), combined with a constant OOP offset field. The skyrmion size is precisely controlled via the applied OOP magnetic field. The magnetic field conditions are varied every 62.5 s (corresponding to 1000 frames) to gather sufficient statistics for analyzing the topological phases. The positions of individual skyrmions are identified using a machine-learning-based, pixel-wise classification algorithm [32]. The detailed analysis procedure can be found in [11]. The experimentally acquired coordinates are analyzed using the same framework as that employed for the simulation data, enabling a direct comparison between experimental and numerical datasets.
2.3  Local Orientational Order Parameter $\psi_{6}(r)$, Correlation Function $c_{6}(r)$, Averaged Absolute Value of Local Orientational Order Parameter $\langle|\Psi_{6}|\rangle$
@@
To identify the topological phases, including liquid, hexatic, and solid phases, the following quantity, termed orientational correlation function, is typically used [5, 11, 12, 13],
@@
n is the number of nearest neighbors of the referenced particle $j$, and $\theta_{m}^{j}$ is the angle between $r_{m}-r_{j}$ and a fixed arbitrary axis. Here, the nearest neighbors are determined using Voronoi tessellation [31]. The factor 6 in the exponent reflects the six-fold rotational symmetry of a perfect hexagonal lattice. When a particle’s six nearest neighbors form an ideal hexagonal arrangement, an absolute value of $\psi_{6}$ of 1 is obtained. The KTHNY theory explains the typical behavior of a correlation function for each topological phase, specifically liquid, hexatic, and solid phases exhibit an exponential decay, algebraic decay, and constant behavior close to 1 as a function of distance $r$, respectively [10, 13].
In addition to identifying the topological phase via the orientational correlation function, another measure of the ordering is the averaged absolute value of the local orientational order parameter $\langle|\Psi_{6}|\rangle$, as described in the following equation [5].

### v4

@@ Abstract
Two-dimensional (2D) particle systems, such as magnetic skyrmions, exhibit topological phase transitions between unique 2D phases. However, a simple and computationally efficient methodology to capture lattice configurational properties and construct an appropriate, easily calculable descriptor for phase identification remains elusive. Here, we propose an indicator for topological phase transitions using persistent homology (PH). PH offers novel insights beyond conventional indicators by capturing topological features derived from the configurational properties of the lattice. The proposed persistent-homology-based indicator, which selectively counts stable features in a persistence diagram, effectively traces the lattice’s ordering changes, as confirmed by comparisons with the conventionally used measure of the ordering (the magnitude of the orientational order parameter $\langle|\Psi_{6}|\rangle$), typically used to identify lattice phases. We demonstrate the applicability of our indicator to experimental data, showing that it yields results consistent with those of simulations. This experimental validation highlights the robustness of the proposed method for real physical systems beyond idealized simulated systems. While our method is demonstrated in the context of skyrmion lattice systems, the approach is general and can be extended to other two-dimensional systems composed of interacting particles. As a key advantage, our indicator offers lower computational complexity than the conventionally used measures.
Introduction
@@
From a fundamental physics perspective, ensembles of skyrmions, known as skyrmion lattices, can exhibit complex phase behavior, including two-dimensional melting transitions [10]. These transitions are theoretically captured by the Kosterlitz–Thouless–Halperin–Nelson–Young (KTHNY) framework, which describes the emergence of intermediate hexatic phases and the role of topological defects such as dislocations and disclinations [10]. Although this behavior has been extensively studied in colloidal systems, skyrmion lattices offer distinct advantages due to their tunability and controllable dynamics that can be fully captured in real time and real space, making them an ideal platform for investigating two-dimensional phase transitions [11].
@@
Despite intensive research, a computationally easy methodology for describing the configurational properties of lattices remains elusive. Conventional methods (e.g., using the local orientational order parameter $\psi_{6}(r)$ [5, 11, 12, 13]) are limited by their reliance on ensemble-averaged quantities and are computationally expensive. That is because they normally average the quantities encoded to each particle over all the particles to identify the state of the systems.
@@
To address this, we introduce a framework based on Topological Data Analysis (TDA), a concept from algebraic topology used to analyze the geometric structure of objects, specifically applying Persistent Homology (PH) to extract configurational properties of lattice configurations [14, 15, 16, 17, 18, 19]. The PH provides a multi-scale view of topological features by analyzing how connected components and loops emerge and disappear across a filtration. This approach has seen successful applications in biological and materials systems [20, 21, 22, 23, 24]. PH can capture and describe the microscopic processes involved in the phase transitions [16]. In this work, we apply PH in a particle-based approach, which models skyrmions as interacting quasi-particles and focuses on their positional configuration that does not require knowledge of the full spin texture. This abstraction enables an efficient computational treatment and highlights the essential geometric features responsible for phase behavior [3, 4, 25]. We propose a topological indicator [26, 27], the Persistent Generator Count with Relative Stability (PGCRS), which selectively counts only the robust topological features of the persistent diagram (PD) in each homology dimension, generated by PH. PGCRS reliably detect phase transitions in the lattice, correlates with conventional measures, and offers a significantly reduced computational complexity. A distinctive aspect of our approach is inverse analysis [28], to trace persistent generators back to specific real-space configurations, enabling a direct physical interpretation of the microscopic structures responsible for topological phase behavior. This conceptual shift from averaging local order to counting persistent topological features offers a new perspective for understanding 2D phase transitions. We apply our framework to experimentally acquired skyrmion configurations, demonstrating the practical applicability of our method for real physical systems. While the focus of this work is on skyrmion systems, the underlying method can be extended to other two-dimensional particle systems.
Results and Discussion
Fig. 1 presents a schematic overview of the analysis workflow. We first (a) obtain the coordinates of skyrmions from the experimental data, and then (b) apply the PH analysis with (c) the calculation of the conventionally used measure of the ordering. Finally, we (d) compare the two indicators. The experimental procedure is detailed in the Methods section. The PH analysis is described in the following subsections and illustrated in Figs. 2–4, where the PDs and their interpretation through inverse analysis are provided. Subsequently, the PH-based indicators and the conventionally used measure are compared in terms of consistency and computational cost, as summarized in Fig. 4 and Table 1. Validation using Molecular Dynamics (MD) simulations is presented in the Supplementary Information (Figs. S1–S6).
Figure 1: Schematic of the analysis workflow: (a) skyrmion coordinates, (b) Persistent Homology (PH) analysis, (c) calculation of a conventional lattice-ordering measure, and (d) comparison between the PH-based indicator and the conventional measure.
Persistent Homology and Persistence Diagrams for Experimental Skyrmion Lattices
Fig. 2 illustrates the filtration process and the corresponding persistence diagrams (PDs) for the 0th- and 1st-degree homology dimensions. The filtration process is performed by continuously increasing the radius of disks centered at the skyrmion coordinates, thereby tracking the emergence and disappearance of topological features such as connected components and loops. The birth and death values in the PDs correspond to the specific filtration parameter $r$ at which these topological features appear and vanish, respectively. The PD0 diagram captures the birth and death of connected components (0th-degree homology), while the PD1 diagram records the birth and death of loop-like structures (1st-degree homology).
In particular, the data points in PD0 indicate when individual skyrmions (disks) merge; thus, all birth values are zero, and the death values correspond to the connection events. The number of data points in PD0 equals the number of skyrmions in the system. In PD1, the lifetime ($\text{death}-\text{birth}$) represents the persistence of loop-like structures formed by connected skyrmions. A large lifetime indicates robust loops that persist over a wide range of the filtration parameter, whereas a small lifetime corresponds to transient loops that quickly disappear as the filtration parameter varies, or possibly to noise.
For a detailed explanation of the PH filtration process, please refer to Ref. [15].
Figure 2: Filtration and persistence diagrams (PDs) in the 0th and 1st homology dimensions. (a) Illustration of the topological features—connected components and loops—emerging during the filtration process, where the radius of disks centered at the skyrmion coordinates is varied. (b1) PD0 (0th-degree homology) and (b2) PD1 (1st-degree homology) with interpretations. For a detailed explanation of the PH filtration process, please refer to Ref. [15].
[…]
In each state, the points labeled (a), (b), and (c) correspond to data points in the PD with a large lifetime (a), small lifetime and small birth value (b), and small lifetime and large birth value (c), respectively. It is observed that data points with a large lifetime originate from a complex structure, while those with a small lifetime originate from a simple structure, consistent with the simulation data. The result reflects the distribution of data points in the PDs (Fig. 3), as the disordered lattices exhibit more complex features compared with ordered lattices. Hence, the PH framework effectively links the actual configurational structure of the skyrmion lattice to the features represented in the corresponding PD, which is difficult to capture with the conventional orientational-order measure.
[…]
Persistent Homology-Based Indicator and Comparison with Conventionally Used Measure of Ordering
Fig. 5 presents the Persistent Generator Count with Relative Stability as a function of $\langle|\Psi_{6}|\rangle$, along with the first derivatives of both indicators. A positive correlation is observed, as confirmed by the Pearson correlation coefficients $r=0.993$ and $r=0.861$ for the values and their derivatives, respectively. The details of the Gaussian Process Regression, used to estimate the first derivatives, are described in Supplementary information and Fig. S7. These results confirm the consistency between the PH-based indicator and the conventional orientational order parameter. As also discussed in the simulation data, the high sensitivity, in particular the first derivatives, supports the validity of the persistent homology-based approach, as phase transitions are characterized by abrupt, non-analytic changes in system properties [30].
Finally, we analyze the computational cost of our indicator compared to the conventionally used indicators. Table 1 shows the comparison in computational complexity for the persistent homology-based indicator and the conventionally used measure $\langle|\Psi_{6}|\rangle$ [12, 31]. It is obvious that the newly constructed indicator has achieved a significant reduction in computational complexity, as the dimension of the system in this work is 2, and therefore the cost is $O(N)$. Furthermore, in the simulation data, the actual runtime for the persistent homology-based indicator is 2.34 seconds per frame, whereas the conventionally used one requires $4.49\times 10^{3}$ seconds in our computational environment.
The ability of persistent homology analysis to capture the ordering may stem from the same geometric foundation as the orientational order parameter, since both rely on Voronoi tessellation, complemented by algebraic topology techniques [32]. One limitation, however, is that the persistent homology-based indicator may require comparisons across different system states to provide meaningful insights. In contrast, the conventionally used measure of the ordering $\langle|\Psi_{6}|\rangle$ has a clear physical interpretation due to its normalization, which ensures values between 0 (disordered state) and 1 (perfect hexagonal order) [5]. Nevertheless, the PH-based indicator can be further generalized through an entropy-based formulation, as algebraic topology provides a well-defined correspondence with thermodynamic quantities [26, 27]. Such a generalization could offer a unified framework for describing both structural complexity and statistical behavior.
Conclusion
In this work, we propose a topological indicator, a Persistent Generator Count with Relative Stability (PGCRS), to characterize phases and phase transitions in two-dimensional quasi-particle systems. As a model system, we use skyrmions, which can be well described as quasi-particles that form lattices in 2D systems. By modeling skyrmions as interacting quasi-particles and applying persistent homology (PH) to their spatial configurations, PGCRS selectively counts stable topological features, providing an interpretable, noise-resistant, and computationally efficient measure of lattice ordering. It correlates with the conventional orientational order parameter $\langle|\Psi_{6}|\rangle$, and reliably traces phase transitions across solid, hexatic, and liquid states. Our inverse analysis reveals that persistent features in the PD with long lifetimes correspond to disordered, complex configurations, while short-lived features are associated with regular, ordered structures. The applicability of our approach is demonstrated using experimental skyrmion lattices, confirming that the PGCRS produces consistent and reliable phase characterization under realistic experimental conditions.
While this work focuses on skyrmion lattice systems, the methodology is broadly applicable to other two-dimensional systems composed of repulsively interacting particles. Future extensions could incorporate insights, such as the Euler entropy, to further strengthen the connection between structural ordering and thermodynamic quantities [26, 27].
Methods
@@
In this work, we treat skyrmions as quasi-particles without considering their detailed spin textures, following previous studies [3, 25, 33]. This abstraction allows us to focus on the configurational properties relevant to phase transitions, while significantly simplifying the computational analysis. We apply PH analysis to experimental skyrmion configurations, while the numerically simulated results are provided in the Supplementary information.
In the subsequent sections, we first describe the experimental setup, and then introduce the persistent homology analysis and present the PH-based indicator. Finally, we describe the conventionally used measure of the ordering.
Experimental Skyrmion Lattices
@@
The experimental skyrmion configuration data used in this work are obtained from magnetic multilayer stacks: Ta(5)/Co20Fe60B20(0.9)/Ta(0.07)/MgO(2)/Ta(5) [3, 11]. Magnetic fields are applied both in-plane (IP) and out-of-plane (OOP) using an Evico Magnetics GmbH Kerr microscope, with the sample maintained at a constant temperature of 333.5 K. Skyrmions are nucleated by applying an IP field pulse under a constant OOP field. The lattice is subsequently equilibrated using an oscillating OOP field at 100 Hz (with amplitudes up to $60\,\upmu\mathrm{T}$), combined with a constant OOP offset field. The skyrmion size is precisely controlled via the applied OOP magnetic field. The magnetic field conditions are varied every 62.5 s (corresponding to 1000 frames) to gather sufficient statistics for analyzing the topological phases. The positions of individual skyrmions are identified using a machine-learning-based, pixel-wise classification algorithm [29]. The detailed analysis procedure can be found in [11].
Persistent Homology for Two-Dimensional Coordinates of Quasi-Particles
In the persistent homology analysis, an alpha-filtration of the 0th- and 1st-degree homology is often applied to the two-dimensional coordinates of quasi-particles [34, 35, 36]. From this, the corresponding persistence diagrams (PD0 and PD1) can be generated [36]. The inverse analysis method of persistent homology is applied to the selected frames of our data, which can identify the original structure corresponding to a specific generator in the persistence diagram [28]. The details of the alpha-filtration, PDs, and inverse analysis are described in Figs. 1–3.
Persistent Homology-Based Indicator
Phase transitions are accompanied by topology changes in the configuration manifold in some classes of systems[16, 26, 27]. We refer the reader to Refs.[26, 27] for a detailed description of the topological invariant and its relationship with phase transitions.:
The topological invariant, Euler characteristic (EC) curve, tracks how the topology of the space changes as the filtration parameter $t$ varies [26], as described in the following equation:
$\chi(t):=\chi(X_{t})=\sum_{k}{(-1)^{k}\beta_{k}(X_{t})},$ (6)
where $X_{t}$ is a topological space, and $\beta_{k}$ is $k$-th Betti number, which counts the number of topological features of dimension $k$. For instance, $\beta_{0}$ and $\beta_{1}$correspond to the number of connected components and loops, respectively. The EC is expressed as the alternating sum of Betti numbers. One important property of the EC is that it is a topological invariant. For example, if two states have different EC values, $\chi_{1}$ and $\chi_{2}$, they are topologically distinct [27]. Thus, by monitoring the evolution of the EC, one can detect topological transitions within the configurational space.
To summarize the EC curve and extract an integer-valued topological invariant, we define the Persistent Generator Count with Relative Stability, as:
where $\beta_{k}^{\text{rel}}(X_{\text{PH}})$ denotes the number of data points in the PD that satisfy specific stability conditions, as described in Fig. 1 and the Supplementary information. In this formulation, the PGCRS provides a single integer summarizing the stable topological features of a single persistence diagram. Although PGCRS is derived from the well-defined topological invariant, the EC, further theoretical justification is required to fully elucidate its implications.
Conventionally Used Measure of Ordering
@@
To identify the topological phases, including liquid, hexatic, and solid phases, the following quantity, termed the orientational correlation function, is typically used [5, 10, 11, 12, 13],
@@
n is the number of nearest neighbors of the referenced particle $j$, and $\theta_{m}^{j}$ is the angle between $r_{m}-r_{j}$ and a fixed arbitrary axis. Here, the nearest neighbors are determined using Voronoi tessellation [37]. The factor 6 in the exponent reflects the sixfold rotational symmetry of a perfect hexagonal lattice. When a particle’s six nearest neighbors form an ideal hexagonal arrangement, an absolute value of $\psi_{6}$ of 1 is obtained. The KTHNY theory explains the typical behavior of a correlation function for each topological phase, specifically, liquid, hexatic, and solid phases exhibit exponential decay, algebraic decay, and constant behavior close to 1 as a function of distance $r$, respectively [10, 13].
In addition to identifying the topological phase via the orientational correlation function, another measure of the ordering is the averaged absolute value of the local orientational order parameter $\langle|\Psi_{6}|\rangle$, as described in the following equation [5]: